OpenStudy (anonymous):

Find the choice below which gives a first step in solving the given system of equations by the addition method 6x - 5y = 12 x + y = 3 Multiply the first equation by 5 and the second by 6. Multiply the first equation by 6 and the second by 5. Multiply the first equation by 5. Multiply the second equation by -6.

OpenStudy (anonymous):

what are your choices?

OpenStudy (anonymous):

Multiply the first equation by 5 and the second by 6. Multiply the first equation by 6 and the second by 5. Multiply the first equation by 5. Multiply the second equation by -6.

OpenStudy (anonymous):

Okay look at those two questions and think what number(s) can I multiply one by so that when I add them together one of the variables cancels

OpenStudy (anonymous):

for instance what number would you multiply the second equation by to cancel out -5y from the first equation?

OpenStudy (anonymous):

6 ?

OpenStudy (anonymous):

lets pull out the y variable from the equations for a second, let the number we would multiply the second equation by in order to cancel out the y in the first equation be A then -5y + ay = 0 y(-5 + a) = 0 -5 + a = 0 a = 5 see how that works now?

OpenStudy (anonymous):

so does a - always turn into a possitive?

OpenStudy (anonymous):

sorry are your equations as follows 6x - 5y = 12 x + y = 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no its because the coefficient of the first y is negative and to cancel out a negative five we add five to both sides of the equation

OpenStudy (anonymous):

6x-5y=12 x+y=3

OpenStudy (anonymous):

Tip for later when writing two different equations press enter between them so that they are on separate lines

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

If the system of equations given below is solved using the substitution method, what is the equation that results from solving the second equation for y, and substituting into the first equation? 3x - 2y = 7 x - y = 5 is the answer y=x-5

OpenStudy (anonymous):

Okay just remember I actually helped you solve for the variable they did not pick for their answer in the multiple choice section (nothing says to only multiply the second equation by five) you will have to use the same process on the x variable but it is a bit of a trick question because you have to add a negative number in order to cancel the x variables in one of the equations

OpenStudy (anonymous):

Another quick tip is to always choose to multiply the equation with the smaller coefficient on the variable you are attempting to cancel out.

OpenStudy (anonymous):

Next time create a new question instead of continuing on in here. Anyways you do get y = x -5 when solving for y using the first equation but you forgot to do the second step, subbing in x - 5 for y in the first equation